# The Born rule – 100 years ago and today

### v2

@@ 2.1 The genesis of the Born rule
The term ’Born rule’ appeared quite late, apparently first in 1934 by Bauer [6, p.302] (”la règle de Born”; cf. Ziegler [76]). Before that, during the gestation period of finding the right level of generalization and interpretation, the pioneers talked more vaguely about Born’s interpretation of quantum mechanics (or of the wave function). For example, in 1927, Jordan [37, p.811] wrote about ”Born’s Deutung der Lösung [der] Schrödingergleichung”.
Wessels [70, p.187] even goes so far as to claim that ”Neither Born nor most of his contemporaries saw in it anything that even suggests the interpretation that we now associate with his name.” Thus it is interesting to consider the genesis of the Born rule, based on the early papers of the pioneers of quantum mechanics. Mehra & Rechenberg [45] give a detailed history of Born’s statistical interpretation, and Bacciagaluppi [5] expands on the early years 1926–1927. However, they miss a number of subtle points that shed new light on the early developments.
@@
The two 1926 papers by Born [8, 9] (the first being a summary of the second) introduced the probabilistic interpretation. His 1926 formulation, ”bestimmt die Wahrscheinlichkeit dafür, daß das aus der $z$-Richtung kommende Elektron in die durch $\alpha,\beta,\gamma$ bestimmte Richtung (und mit einer Phasenänderung $\delta$) geworfen wird” (Born [8, p.865f]) when rephrased in modern terminology, is a special case of the objective scattering form (BR-OSc) of the Born rule. Note that Born didn’t have the concept of an S-matrix, first introduced in 1937 by Wheeler [72]. However, the square of the absolute value the S-matrix elements for single-particle scattering is proportional111 The S-matrix elements $S_{ik}$ are the inner products $e_{i}^{*}Se_{k}$, where the $e_{i}$ are the standard basis vectors of the matrix $S$; in Born’s case the stationary states of the particle. The relation with the S-matrix is visible more explicitly in Dirac [20], who computed in 1927 what are today called the S-matrix elements in the so-called Born approximation, to the ”Ausbeutefunktion” $\Phi_{n,m}$ mentioned in Born [9, p.824], already used (with similar notation) in Nordheim [60, p.512].
@@
The 1927 paper by Born [11] extends this rule on p.173 to probabilities for quantum jumps (”Quantensprung”, p.172) between energy eigenstates, given by the absolute squares of inner products of the corresponding eigenstates, still using objective rather than measurement-based language: For a system initially in state $n$ given by formula (9) in Born’s paper he says, ”Das Quadrat $|b_{nm}|^{2}$ ist gemäß unserer Grundhypothese die Wahrscheinlichkeit dafür, daß das System sich nach Ablauf der Störung im Zustand $m$ befindet” Here state $n$ is the $n$th stationary state (eigenstate with a time-dependent harmonic phase) of the Hamiltonian.
Born derives this rule from two assumptions. The first assumption, made on p.170 and repeated on p.171 after (5), is that an atomic system is always in a definite stationary state: ”Wir werden also an dem Bohrschen Bilde festhalten, daß ein atomares System stets nur in einem stationaren Zustand ist. […] im allgemeinen aber werden wir in einem Augenblick nur wissen, daß auf Grund der Vorgeschichte und der bestehenden physikalischen Bedingungen eine gewisse Wahrscheinlichkeit dafür besteht, daß das Atom im $n$-ten Zustand ist.”
@@
p.4: ”Im eigentlichen Gebiet der neuen Dynamik kann man sich meistens auf das Prinzip verlassen, dass das Quadrat der Wellenamplitude, die Intensität, an einem bestimmten Ort und zu einer bestimmten Zeit die Wahrscheinlichkeit dafür liefert, dass sich das Teilchen zu dieser Zeit an diesem Ort befindet. Man überlegt leicht, dass dieses Prinzip notwendig ist, um Beugung und Interferenz zu erklären.”
@@
p.6f, after mentioning difficulties of the objective position-based view: ”Schliesslich gibt es noch eine vierte Betrachtungsweise, die augenblicklich die meisten Anhänger hat. Sie wurde von Heisenberg und Bohr entwickelt. […] Nach dieser Auffassung ist die Welle […] nur eine symbolisch Darstellung von dem, was wir über das Teilchen wissen. […] so gestatten Intensitätsverteilung und Spektralbeschaffenheit der Welle, die Wahrscheinlichkeit dafür anzugeben, dass ein zur Zeit $t$ ausgeführtes Experiment das Teilchen an einem bestimmten Ort findet oder ihm einen bestimmten Bewegungszustand zuschreibt.”
@@
The objective interpretation is needed to be able to argue on p.70 as follows: ”Die Dichte des Schwarms kann man auch als die Wahrscheinlichkeit dafür betrachten, dass ein Teilchen, dessen Bewegung zur betrachteten Klasse gehört, sich zu einer bestimmten Zeit an einem bestimmten Punkt befindet, wenn seine genaue Lage unbekannt ist. […] Wir hatten diese Annahme in der Einleitung als ’Interferenzprinzip’ bezeichnet.” However, de Broglie mentions at several places the associated difficulties.
@@
In the main text, he mentions on p.117f einerseits $a^{2}=|\psi(x)|^{2}$ (”nach dem Interferenzprinzip” of Schrödinger, pp.3–5), and andrerseits $a_{k}^{2}$ (”nach dem Bornschen Postulat”), now in an objective formulation: ”Als die Wellenmechanik noch am Anfang ihrer Entwicklung stand, hat Max Born schon den Vorschlag gemacht, jede Grösse $a_{k}^{2}$ als die relative Wahrscheinlichkeit dafür aufzufassen, dass das Teilchen den zu $\psi$ gehörigen Bewegungszustand hat. […] Wir nennen dieses Bornsche Postulat das ’Prinzip der spektralen Zerlegung’. Nimmt man seine Gültigkeit an, so hat die Bestimmung des Teilchens durch die zugehörige Welle eine doppelte Unbestimmtheit: einerseits ist die Lage des Teilchens unbestimmt, da es nach dem Interferenzprinzip eine endliche Wahrscheinlichkeit gibt, dass man das Teilchen an irgendeiner Stelle des Raumgebiets findet, das vom Wellenzug ausgefüllt wird, und zwar ist sie gleich der resultierenden Intensität $a^{2}$. Andererseits ist der durch Impuls und Energie gemessene Bewegungszustand nach dem Prinzip der spektralen Zerlegung ebenfalls unbestimmt, weil es mehrere mögliche Bewegungszustände gibt und die Wahrscheinlichkeit für jeden durch das Amplitudenquadrat der zugehörigen monochromatischen Komponente in der spektralen Zerlegung des Wellenzugs ist.”
@@
De Broglie uses here exactly the same terminology as Born did in his 1927 paper [10, p.168], ”Die Teilchen sind immer von einem Wellenvorgang begleitet; diese de Broglie–Schrödingerschen Wellen hängen von den Kräften ab und bestimmen durch das Quadrat ihrer Amplitude die Wahrscheinlichkeit dafür, dass ein Teilchen in dem der Welle entsprechenden Bewegungszustand vorhanden ist. In seiner letzten Mitteilung beschäftigt sich auch Schrödinger mit dem Amplitudenquadrat der Wellen und führt dafür den Ausdruck ’Gewichtsfunktion’ ein, der sich der Terminologie der Statistik schon sehr nähert.” Thus we know that ”Bewegungszustand” stands for the energy-momentum eigenstate of the particle. Indeed, ”Bewegungszustand” (”state of motion”) was at that time a technical term now no longer in use. The 1898 book by Jordan [36, p.1f] defines: ”Mit dem Worte Kraft bezeichnet man die Ursache einer Änderung des Bewegungszustandes eines Körpers. – Als Bewegungszustand ist nicht nur eine Bewegung irgend welcher Art, sondern auch die Ruhe – als Abwesenheit jeglicher Bewegung (Geschwindigkeit = 0, vgl. S. 3) – aufzufassen. […] Jeder Körper behält den Bewegungszustand, den er in irgend einem Momente hat, nach Richtung und Geschwindigkeit unverändert bei, so lange keine äussere Kraft (ändernd) auf ihn einwirkt.” Thus Bewegungszustand is a state that persists unless changed by an interaction.
@@
In 1925, Kramers & Heisenberg [39, p.692] use the term in the quantum context for stationary states: ”die sich auf zwei Bewegungszustände beziehen, für die die Werte der Grössen $J_{1}\cdots J_{s}$ sich um $\tau_{1}h\cdots\tau_{s}h$ unterscheiden. […] die Bewegungszustände seien durch Punkte in einer $J_{1},J_{2}$-Ebene, der Zeichenebene der Figur, dargestellt.”
@@ 2.2 Early objective, measurement-independent formulations
(B1) A state of a quantum system at time $t$ is a stationary state,222 Bacciagaluppi [5] comments this in his footnote 9 as follows: ”This is not in fact stated explicitly but has to be read between the lines.” The preceding subsection may thus be viewed as giving details on how to read between the lines. given by a simultaneous eigenstate of $H$, ${\bf p}$, and $J^{2}$.
@@
Born’s interpretation met with strong resistance by Schrödinger, who allegedly said – as Heisenberg [33, p.6] recalls much later – at a conference in Kopenhagen in September 1926, ”Wenn es doch bei dieser verdammten Quantenspringerei bleiben soll, dann bedauere ich, dass ich mich überhaupt mit diesem Gegenstand beschäftigthabe.”. An exchange of letters between him and Born displays the two extreme poles between which the quantum community had to find a consensus in the subsequent years.
@@
Letter of Schrödinger to Born (November 2, 1926) (von Meyenn [46, p.329]): ”Ich habe eben Ihre letzte, ebenerhaltene Arbeit über den Adiabatensatz überflogen (also noch nicht wirklich gelesen). […] Ich habe aber doch den Eindruck, dass Sie und andere, die im Wesentlichen Ihre Ansicht teilen, zu tief im Banne derjenigen Begriffe stehen (wie stationäre Zustände, Quantensprünge usw.), die sich in den letzten zwölf Jahren Bürgerrecht in unserem Denken erworben haben, um dem Versuch, aus diesem Denkschema wieder herauszukommen, volle Gerechtigkeit widerfahren zu lassen. Als Beispiel führe ich z. B. an, dass Sie zur Kritik der Annahme: im Atom seien mehrere Eigenschwingungen gleichzeitig erregt, die Frage stellen: hat es Sinn, zu sagen, das Atom befinde sich gleichzeitig in mehreren stationären Zuständen? Ganz ferne scheint Ihnen die andere naheliegende Bemerkung zu liegen: immer und überall sonst, wo wir mit schwingungsfähigen Systemen zu tun haben, schwingen dieselben im Allgemeinen nicht in der Form einer Eigenschwingung, sondern mit einem Gemisch derselben.”
@@
Letter of Born to Schrödinger (November 6, 1926): (von Meyenn [46, p.333f]): ”Als direkteste Evidenz der stationären Zustände haben wir die Elektronenstösse, und hier denke ich vor allem an die Stösse zweiter Art. Wir können wohl als Tatsache nehmen, dass ein Elektron beim Zusammenstoss mit einem Atom, das in irgend einem Zustande ist, immer nur einen Energiezuwachs bekommt. Wenn nun in einem Atom mehrere Eigenschwingungen gleichzeitig angeregt sein könnten, wie sollte man das verstehen? Was nützt da Ihr Gegenargument, dass das Elektron selber eine ’Wellengruppe’ sein könne? Es handelt sich doch um die Atomwellen, die diskreten Stufen entsprechen. Natürlich ist es möglich zu sagen: das in mehreren Frequenzen schwingende Atom reagiert eben in besonderer Weise so, dass es nur die einer Frequenz entsprechende Energie auf das Elektron überträgt. Aber das ist doch nur eine Ausrede. Mir scheint, dass die Bohrsche Redeweise die natürliche Beschreibung und Zusammenfassung eines grossen Tatsachenbereichs ist; darum ist es Aufgabe jeder feineren Theorie, diese Redeweise zu rechtfertigen.”
@@
Letter of Pauli to Heisenberg, Oktober 19, 1926 (Pauli [62, p.347]): ”Nun ist es so: alle Diagonalelemente der Matrizen (wenigstens von Funktionen der $p$ allein oder der $q$ allein) kann man überhaupt schon jetzt kinematisch deuten. Denn man kann ja zunächst nach der Wahrscheinlichkeit fragen, dass in einem bestimmten stationären Zustand des Systems die Koordinaten $q_{k}$ seiner Teilchen ($k=1,\ldots,f$) zwischen $q_{k}$ und $q_{k}+dq_{k}$ liegen. Die Antwort hierauf ist $|\psi(q_{1},\ldots,q_{f})|^{2}dq_{1}\ldots dq_{f}$, wenn $\psi$ die Schrödingersche Eigenfunktion ist.”
The position probability density interpretation was published in the next year in a footnote of Pauli [61, p.83,Footnote 1]: ”Wir wollen diese […] Funktion im Sinne der von Born in seiner Stoßmechanik [here he cites [8, 9]] vertretenen Auffassung des ”Gespensterfeldes” folgendermaßen deuten: Es ist $|\psi(q_{1}\ldots q_{f})|^{2}dq_{1}\ldots dq_{f}$ die Wahrscheinlichkeit dafür, daß im betreffenden Quantenzustand des Systems diese Koordinaten sich zugleich im betreffenden Volumenelement $dq_{1}\ldots dq_{f}$ des Lageraums befinden.” Apart from its objective formulation (no reference to measurement), this is a special case of the universal formulation (BR-US) of the Born rule.
The 1927 paper by von Neumann [55, p.45] generalized Pauli’s statement to arbitrary selfadjoint operators, again stated as an objective (i.e., measurement independent) interpretation. For discrete energy spectra and their energy levels, we still read p.48: ”unquantized states are impossible” (”nicht gequantelte Zustände sind unmöglich”). Thus he follows Born’s view that all physically possible states are stationary.
@@
The first published statement of (BR-OE) seems to be the 1927 paper by Landau [40]. The formulas (4a) and (5), corresponding in modern notation to $\langle A\rangle:={\mathop{\rm Tr}\,}\rho A$ in the mixed case and $\langle A\rangle:=\psi^{*}A\psi$ in the pure case, $\rho=\psi\psi^{*}$, are interpreted in Footnote 2 as probability mean (”Wahrscheinlichkeitsmittelwert”). Though he mentions that this rule is already known, earlier formulas in print for the mean always assumed a representation in which the operator is already diagonal. Again there is no reference to measurement.

### v3

@@ 2.1 The genesis of the Born rule
The term ’Born rule’ appeared quite late, apparently first in 1934 by Bauer [7, p.302] (”la règle de Born”; cf. Ziegler [81]). Before that, during the gestation period of finding the right level of generalization and interpretation, the pioneers talked more vaguely about Born’s interpretation of quantum mechanics (or of the wave function). For example, in 1927, Jordan [41, p.811] wrote about ”Born’s interpretation of the solution [of the] Schrödinger equation”
(German original: ”Born’s Deutung der Lösung [der] Schrödingergleichung”).
Wessels [75, p.187] even goes so far as to claim that ”Neither Born nor most of his contemporaries saw in it anything that even suggests the interpretation that we now associate with his name.” Thus it is interesting to consider the genesis of the Born rule, based on the early papers of the pioneers of quantum mechanics. Mehra & Rechenberg [49] give a detailed history of Born’s statistical interpretation, and Bacciagaluppi [5] expands on the early years 1926–1927. However, they miss a number of subtle points that shed new light on the early developments.
@@
The two 1926 papers by Born [9, 10] (the first being a summary of the second) introduced the probabilistic interpretation. His 1926 formulation ”determines the probability that the electron cominf from the $z$-direction is thrown into the direction determined by $\alpha,\beta,\gamma$ (with a phase change of $\delta$)”
(German original, Born [9, p.865f]: ”bestimmt die Wahrscheinlichkeit dafür, dass das aus der $z$-Richtung kommende Elektron in die durch $\alpha,\beta,\gamma$ bestimmte Richtung (und mit einer Phasenänderung $\delta$) geworfen wird”) when rephrased in modern terminology, is a special case of the objective scattering form (BR-OSc) of the Born rule. Note that Born didn’t have the concept of an S-matrix, first introduced in 1937 by Wheeler [77]. (The S-matrix elements $S_{ik}$ are the inner products $e_{i}^{*}Se_{k}$, where the $e_{i}$ are the standard basis vectors of the matrix $S$; in Born’s case the stationary states of the particle. The relation with the S-matrix is visible more explicitly in Dirac [23], who computed in 1927 what are today called the S-matrix elements in the so-called Born approximation.) However, the square of the absolute value the S-matrix elements for single-particle scattering is proportional to the ”Ausbeutefunktion” $\Phi_{n,m}$ mentioned in Born [10, p.824], already used (with similar notation) in Nordheim [64, p.512].
@@
The 1927 paper by Born [12] extends this rule on p.173 to probabilities for quantum jumps (”Quantensprung”, p.172) between energy eigenstates, given by the absolute squares of inner products of the corresponding eigenstates, still using objective rather than measurement-based language: For a system initially in state $n$ given by formula (9) in Born’s paper he says, ”The square $|b_{nm}|^{2}$ is, according to our basic hypothesis, the probability that after the end of the disturbance, the system is in state $m$
(German original: ”Das Quadrat $|b_{nm}|^{2}$ ist gemäss unserer Grundhypothese die Wahrscheinlichkeit dafür, dass das System sich nach Ablauf der Störung im Zustand $m$ befindet”) Here state $n$ is the $n$th stationary state (eigenstate with a time-dependent harmonic phase) of the Hamiltonian.
Born derives this rule from two assumptions. The first assumption, made on p.170 and repeated on p.171 after (5), is that an atomic system is always in a definite stationary state: ”Thus we shall stick to Bohr’s picture that an atomic system is always only in a stationary state. […] but in general we will know in any moment only that, because of the prior history and the given physical conditions, there is a certain probability that the atom is in the $n$th state.”
(German original: ”Wir werden also an dem Bohrschen Bilde festhalten, dass ein atomares System stets nur in einem stationaren Zustand ist. […] im allgemeinen aber werden wir in einem Augenblick nur wissen, dass auf Grund der Vorgeschichte und der bestehenden physikalischen Bedingungen eine gewisse Wahrscheinlichkeit dafür besteht, dass das Atom im $n$-ten Zustand ist.”)
@@
[…]
p.4: ”In the proper domain of the new dynamics, one can usually rely on the principle that the square of the wave amplitude, the intensity, at a particular place and a particular time gives the probability that the particle is at this time at this place. One ckecks easily that this principle is necessary to explain diffraction and interference.”
@@
(German original: ”Im eigentlichen Gebiet der neuen Dynamik kann man sich meistens auf das Prinzip verlassen, dass das Quadrat der Wellenamplitude, die Intensität, an einem bestimmten Ort und zu einer bestimmten Zeit die Wahrscheinlichkeit dafür liefert, dass sich das Teilchen zu dieser Zeit an diesem Ort befindet. Man überlegt leicht, dass dieses Prinzip notwendig ist, um Beugung und Interferenz zu erklären.”)
p.6f, after mentioning difficulties of the objective position-based view: ”Finally, there is still a fourth approach, which currently has the most supporters. It was developed by Heisenberg and Bohr. […] According to this view , the wave […] is only a symbolic representation of what we know about the particle. […] Thus the intensity distribution and spectral nature of the wave allws one to give the probability that an experiment performed at time $t$ finds the particle at a particular place, or assigns to it a particular state of motion.”
@@
(German original: ”Schliesslich gibt es noch eine vierte Betrachtungsweise, die augenblicklich die meisten Anhänger hat. Sie wurde von Heisenberg und Bohr entwickelt. […] Nach dieser Auffassung ist die Welle […] nur eine symbolisch Darstellung von dem, was wir über das Teilchen wissen. […] so gestatten Intensitätsverteilung und Spektralbeschaffenheit der Welle, die Wahrscheinlichkeit dafür anzugeben, dass ein zur Zeit $t$ ausgeführtes Experiment das Teilchen an einem bestimmten Ort findet oder ihm einen bestimmten Bewegungszustand zuschreibt.”)
The objective interpretation is needed to be able to argue on p.70 as follows: ”The density of the ensemble can also be regarded as the probability that a particle, whose motion belongs to the class considered, is at a particular time at a particular point, if its exact position is unknown. […] We had denoted this assumption in the introduction as ’interference principle’.”
@@
(German original: ”Die Dichte des Schwarms kann man auch als die Wahrscheinlichkeit dafür betrachten, dass ein Teilchen, dessen Bewegung zur betrachteten Klasse gehört, sich zu einer bestimmten Zeit an einem bestimmten Punkt befindet, wenn seine genaue Lage unbekannt ist. […] Wir hatten diese Annahme in der Einleitung als ’Interferenzprinzip’ bezeichnet.”) However, de Broglie mentions at several places the associated difficulties.
In the main text, he mentions on p.117f on the one hand $a^{2}=|\psi(x)|^{2}$ (”by the interference principle” of Schrödinger, pp.3–5), and on the other hand $a_{k}^{2}$ (”by the postulate of Born”), now in an objective formulation: ”When the wave mechanics was still at the beginning of its development, Max Born already made the proposal to regard every quantity $a_{k}^{2}$ as the relative probability for a particle to have the state of motion belonging to $\psi$. […] We call this postulate of Born the ’principle of spectral resolution’. Assuming its validity, the determination of the particle through the corresponding wave has a double uncertainty: On the one hand, the position of the particle is undetermined, since by the interference principle, there is a positive probability that one finds the particle in any place of the spatial region occupied by the wave packet, being equal to the resulting intensity $a^{2}$. On the other hand, the state of motion, measured through momentum and energy, is by the principle of spectral resolution also undetermined, since there are several possible states of motion, and the probability for each one is the amplitude square of the corresponding monochromatic component in the spectral resolution of the wave packet.”
@@
(German original: ”Als die Wellenmechanik noch am Anfang ihrer Entwicklung stand, hat Max Born schon den Vorschlag gemacht, jede Grösse $a_{k}^{2}$ als die relative Wahrscheinlichkeit dafür aufzufassen, dass das Teilchen den zu $\psi$ gehörigen Bewegungszustand hat. […] Wir nennen dieses Bornsche Postulat das ’Prinzip der spektralen Zerlegung’. Nimmt man seine Gültigkeit an, so hat die Bestimmung des Teilchens durch die zugehörige Welle eine doppelte Unbestimmtheit: einerseits ist die Lage des Teilchens unbestimmt, da es nach dem Interferenzprinzip eine endliche Wahrscheinlichkeit gibt, dass man das Teilchen an irgendeiner Stelle des Raumgebiets findet, das vom Wellenzug ausgefüllt wird, und zwar ist sie gleich der resultierenden Intensität $a^{2}$. Andererseits ist der durch Impuls und Energie gemessene Bewegungszustand nach dem Prinzip der spektralen Zerlegung ebenfalls unbestimmt, weil es mehrere mögliche Bewegungszustände gibt und die Wahrscheinlichkeit für jeden durch das Amplitudenquadrat der zugehörigen monochromatischen Komponente in der spektralen Zerlegung des Wellenzugs ist.”)
De Broglie uses here exactly the same terminology as Born did in his 1927 paper [11, p.168], ”The particles are always accompanied by a wave process; these de Broglie–Schrödinger waves depend on the forces and determine through the square of their amplitude the probability that a particle is present in the state of motion corresponding to the wave. In hist last communication, Schrödinger also discussed the amplitude square of the waves and introduces for it the term ’weight function’, which is already very close to the terminology of statistics. ”
@@
(German original: ”Die Teilchen sind immer von einem Wellenvorgang begleitet; diese de Broglie–Schrödingerschen Wellen hängen von den Kräften ab und bestimmen durch das Quadrat ihrer Amplitude die Wahrscheinlichkeit dafür, dass ein Teilchen in dem der Welle entsprechenden Bewegungszustand vorhanden ist. In seiner letzten Mitteilung beschäftigt sich auch Schrödinger mit dem Amplitudenquadrat der Wellen und führt dafür den Ausdruck ’Gewichtsfunktion’ ein, der sich der Terminologie der Statistik schon sehr nähert.”) Thus we know that ”Bewegungszustand” (”state of motion”) stands for the energy-momentum eigenstate of the particle. Indeed, ”state of motion” was at that time a technical term now no longer in use. The 1898 book by Jordan [40, p.1f] defines: ”With the word ’force’, one denotes the cause of a change of the state of motion of a body. – As state of motion one has to understand not only a motion of any kind but also the rest – as lack of any motion (velocity = 0, cf. p.3). […] Every body maintains without change the state of motion that it has at any motion, in direction and speed, as long as no external force affects a change.”
(German original: ”Mit dem Worte Kraft bezeichnet man die Ursache einer Änderung des Bewegungszustandes eines Körpers. – Als Bewegungszustand ist nicht nur eine Bewegung irgend welcher Art, sondern auch die Ruhe – als Abwesenheit jeglicher Bewegung (Geschwindigkeit = 0, vgl. S. 3) – aufzufassen. […] Jeder Körper behält den Bewegungszustand, den er in irgend einem Momente hat, nach Richtung und Geschwindigkeit unverändert bei, so lange keine äussere Kraft (ändernd) auf ihn einwirkt.”) Thus Bewegungszustand is a state that persists unless changed by an interaction. In 1925, Kramers & Heisenberg [43, p.692] use the term in the quantum context for stationary states: ”which refer to two states of motion for which the values of the quantities $J_{1}\cdots J_{s}$ differ by $\tau_{1}h\cdots\tau_{s}h$. […] We represent the states of motion by points in a $J_{1},J_{2}$-plane, the drawing plane of the figure.”
@@
(German original: ”die sich auf zwei Bewegungszustände beziehen, für die die Werte der Grössen $J_{1}\cdots J_{s}$ sich um $\tau_{1}h\cdots\tau_{s}h$ unterscheiden. […] die Bewegungszustände seien durch Punkte in einer $J_{1},J_{2}$-Ebene, der Zeichenebene der Figur, dargestellt.”)
@@ 2.2 Early objective, measurement-independent formulations
(B1) A state of a quantum system at time $t$ is a stationary state, given by a simultaneous eigenstate of $H$, ${\bf p}$, and $J^{2}$. (Bacciagaluppi [5] comments this in his footnote 9 as follows: ”This is not in fact stated explicitly but has to be read between the lines.” The preceding subsection may thus be viewed as giving details on how to read between the lines.)
@@
Born’s interpretation met with strong resistance by Schrödinger, who allegedly said – as Heisenberg [37, p.6] recalls much later – at a conference in Kopenhagen in September 1926, ”But should this damned quantum jumping persist, then I regret having even worked on this subject.”
@@
(German original: ”Wenn es doch bei dieser verdammten Quantenspringerei bleiben soll, dann bedauere ich, dass ich mich überhaupt mit diesem Gegenstand beschäftigt habe.”.) An exchange of letters between him and Born displays the two extreme poles between which the quantum community had to find a consensus in the subsequent years.
Letter of Schrödinger to Born (November 2, 1926) (von Meyenn [50, p.329]): ”I just skimmed (hence not yet really read) your last, just received, paper on the adiabatic theorem. […] But I still have the impression that you, and others who essentially share your view, are too deep under the spell of those concepts (such as stationary states, quantum jumps, etc.) that in the last twelve years established their right to be present in our thought, to give full justice to the attempt to escape again form this thought pattern. As an example I mention, e.g., that to the critique of the assumption that in the atom several eigen oscillations are simultaneously excited, you ask the question, does it make sense to say the atom is simultaneously in several eigen oscillations? Aother remark seems to be to you a long way off: Always and everywhere else, where we have to do with oscillating systems, these generally oscillate not in the form of an eigen oscillation but in a superposition of these.”
@@
(German original: ”Ich habe eben Ihre letzte, ebenerhaltene Arbeit über den Adiabatensatz überflogen (also noch nicht wirklich gelesen). […] Ich habe aber doch den Eindruck, dass Sie und andere, die im Wesentlichen Ihre Ansicht teilen, zu tief im Banne derjenigen Begriffe stehen (wie stationäre Zustände, Quantensprünge usw.), die sich in den letzten zwölf Jahren Bürgerrecht in unserem Denken erworben haben, um dem Versuch, aus diesem Denkschema wieder herauszukommen, volle Gerechtigkeit widerfahren zu lassen. Als Beispiel führe ich z. B. an, dass Sie zur Kritik der Annahme: im Atom seien mehrere Eigenschwingungen gleichzeitig erregt, die Frage stellen: hat es Sinn, zu sagen, das Atom befinde sich gleichzeitig in mehreren stationären Zuständen? Ganz ferne scheint Ihnen die andere naheliegende Bemerkung zu liegen: immer und überall sonst, wo wir mit schwingungsfähigen Systemen zu tun haben, schwingen dieselben im Allgemeinen nicht in der Form einer Eigenschwingung, sondern mit einem Gemisch derselben.”)
Letter of Born to Schrödinger (November 6, 1926): (von Meyenn [50, p.333f]): ”As most direct evidence of the stationary states we have the electron collisions, and here I think primarily of collisions of the second kind. We may taske, I think, as fact that an electron colliding with an atom in some particular state always gets only one increase in energy. Suppose now that in an atom several eigen oscillations can be simultaneously excited, how should one understand this? What is the use of your conter argument that the electron could itself be a ’wave group’? For it is the atomic waves that correspond to discrete steps. Of course it is possible to say: The aton that oscillates in several frequencies responds in a special way, such that it transmits to the electron only the energy corresponding to one frequency. But this is only an excuse. It seems to me that Bohr’s way of speaking is the natural description and summary of a large body of facts; therefore it is the task of every finer theory to justify this way of speaking.”
@@
(German original: ”Als direkteste Evidenz der stationären Zustände haben wir die Elektronenstösse, und hier denke ich vor allem an die Stösse zweiter Art. Wir können wohl als Tatsache nehmen, dass ein Elektron beim Zusammenstoss mit einem Atom, das in irgend einem Zustande ist, immer nur einen Energiezuwachs bekommt. Wenn nun in einem Atom mehrere Eigenschwingungen gleichzeitig angeregt sein könnten, wie sollte man das verstehen? Was nützt da Ihr Gegenargument, dass das Elektron selber eine ’Wellengruppe’ sein könne? Es handelt sich doch um die Atomwellen, die diskreten Stufen entsprechen. Natürlich ist es möglich zu sagen: das in mehreren Frequenzen schwingende Atom reagiert eben in besonderer Weise so, dass es nur die einer Frequenz entsprechende Energie auf das Elektron überträgt. Aber das ist doch nur eine Ausrede. Mir scheint, dass die Bohrsche Redeweise die natürliche Beschreibung und Zusammenfassung eines grossen Tatsachenbereichs ist; darum ist es Aufgabe jeder feineren Theorie, diese Redeweise zu rechtfertigen.”)
@@
Letter of Pauli to Heisenberg, Oktober 19, 1926 (Pauli [67, p.347]): ”Now it is thus: all diagonal elements of the matrices (at least of functions of $p$ alone or of $q$ alone) one can interpret kinematically already now. For one can first ask for the probability that in a particular stationary state, the coordinates $q_{k}$ of the particles ($k=1,\ldots,f$) lie between $q_{k}$ und $q_{k}+dq_{k}$. The answer is $|\psi(q_{1},\ldots,q_{f})|^{2}dq_{1}\ldots dq_{f}$, if $\psi$ is Schrödinger’s eigenfunction.”
(German original: ”Nun ist es so: alle Diagonalelemente der Matrizen (wenigstens von Funktionen der $p$ allein oder der $q$ allein) kann man überhaupt schon jetzt kinematisch deuten. Denn man kann ja zunächst nach der Wahrscheinlichkeit fragen, dass in einem bestimmten stationären Zustand des Systems die Koordinaten $q_{k}$ seiner Teilchen ($k=1,\ldots,f$) zwischen $q_{k}$ und $q_{k}+dq_{k}$ liegen. Die Antwort hierauf ist $|\psi(q_{1},\ldots,q_{f})|^{2}dq_{1}\ldots dq_{f}$, wenn $\psi$ die Schrödingersche Eigenfunktion ist.”)
The position probability density interpretation was published in the next year in a footnote of Pauli [65, p.83,Footnote 1]: ”We want to interpret this […] function in the spirit of the ghost field view of Born given in his book [9, 10] as follows: $|\psi(q_{1}\ldots q_{f})|^{2}dq_{1}\ldots dq_{f}$ is the probability that in the respective quantum state of the system these coordinates lie simultaneously in the respective volume element $dq_{1}\ldots dq_{f}$ of position space.”
(German original: ”Wir wollen diese […] Funktion im Sinne der von Born in seiner Stossmechanik [[9, 10]] vertretenen Auffassung des ”Gespensterfeldes” folgendermassen deuten: Es ist $|\psi(q_{1}\ldots q_{f})|^{2}dq_{1}\ldots dq_{f}$ die Wahrscheinlichkeit dafür, dass im betreffenden Quantenzustand des Systems diese Koordinaten sich zugleich im betreffenden Volumenelement $dq_{1}\ldots dq_{f}$ des Lageraums befinden.”) Apart from its objective formulation (no reference to measurement), this is a special case of the universal formulation (BR-US) of the Born rule.
The 1927 paper by von Neumann [59, p.45] generalized Pauli’s statement to arbitrary selfadjoint operators, again stated as an objective (i.e., measurement independent) interpretation. For discrete energy spectra and their energy levels, we still read p.48: ”unquantized states are impossible”
(German original: ”nicht gequantelte Zustände sind unmöglich”). Thus he follows Born’s view that all physically possible states are stationary.
@@
The first published statement of (BR-OE) seems to be the 1927 paper by Landau [44]. The formulas (4a) and (5), corresponding in modern notation to $\langle A\rangle:={\mathop{\rm tr}\,}\rho A$ in the mixed case and $\langle A\rangle:=\psi^{*}A\psi$ in the pure case, $\rho=\psi\psi^{*}$, are interpreted in Footnote 2 as probability mean (German original: ”Wahrscheinlichkeitsmittelwert”). Though he mentions that this rule is already known, earlier formulas in print for the mean always assumed a representation in which the operator is already diagonal. Again there is no reference to measurement.